# On Topology and Geometry of Hessian Manifolds

### v1

@@ 6.3 The obstruction line bundle
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type, and $(\mathscr{L},D)$ its obstruction line bundle. Then there exists a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}\rightarrow M$ such that the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is a 1-form potential of $((M,\nabla),h)$. Moreover, if we assume further that there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$, then $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle.
@@
7 Digression on Complex Geometry
The ultimate objective of this digression section is to prove that every positive flat line bundles over an affine manifold $(M,\nabla)$ can be lifted to a positive holomorphic line bundles over the Sasaki lifting of $(M,\nabla)$.
We shall first recall the concept of Sasaki lifting from complex analytic geometry.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon TM\rightarrow M$ be the tangent bundle of $M$. Denote by $\mathcal{A}$ the sheaf of affine functions on $(M,\nabla)$. Consider any coordinate map
@@ 7 Digression on Complex Geometry
of $M$ such that $x^{1},\dots,x^{n}\in\mathcal{A}(U)$ are affine functions, and that $TM\rightarrow M$ is trivialized over $U$. Then
@@
is a coordinate map of $TM$. Now, the canonical almost complex structure $J_{\nabla}$ on the tangent bundle $T(TM)\rightarrow TM$ is defined by
@@
for all $i=1,\dots,n$. Then by [10] the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of the affine manifold $(M,\nabla)$.
@@
8 Legendre Transformation
It turns out that there exists a duality between radiant Hessian manifolds and Hessian manifolds of Koszul type, resembling the classical Legendre transformation. In this section, we study Hessian manifold of Koszul type in terms of this generalized Legendre transform.
@@ 8 Legendre Transformation
It is readily seen that, if $H_{1}$ and $H_{2}$ are vector fields on an affine manifold such that $H_{1}$ is Euler, then $H_{2}$ is also Euler if and only if $H_{1}-H_{2}$ is parallel.
@@
For simplicity, we identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
@@
on $(\Omega,\pi^{*}\nabla)$, where $a^{1},\dots,a^{n}\in\mathbb{R}$. Moreover, by Poincaré-Hopf theorem, we have that $(a^{1},\dots,a^{n})\notin\Omega$. Up to an affine change of coordinates, we may assume $a^{1},\dots,a^{n}=0$. Since $M$ is compact, it follows that $H$ is a complete vector field. Since $\pi\colon\Omega\rightarrow M$ is a covering space, we obtain that $\pi^{*}H$ is also a complete vector field as $H$ is. Now it remains to recall that the integral curves of the standard Euler vector field
$x^{1}\cdot\partial/\partial x^{1}+\cdots+x^{n}\cdot\partial/\partial x^{n}$
@@
Corollary 8.6.
Let $(M,\nabla)$ be a compact radiant affine manifold. Then $M$ is aspherical.
Proof.
This follows from Proposition8.5, and the fact that a cone is a deformation retraction to its apex. ∎
@@
then the constructions in Theorem8.10 and Theorem8.11 together recover the Legendre transformation in classical mechanics.
9 Space-like Lagrangian Immersions
@@
For an affine manifold $(M,\nabla)$, the pseudo-Riemannian metric
$g:=\frac{1}{2}(dx^{i}\otimes dp_{i}+dp_{i}\otimes dx^{i})$
on $T^{*}M$ is called the pseudo-Riemannian metric induced by $\nabla$, where $x^{1},\dots,x^{n}$ are affine coordinates of $M$ and $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ are canonical coordinates on $T^{*}M$.
Since the second derivatives of affine functions vanish, the notion of induced pseudo-Riemannian metrics is well-defined. Note that the induced pseudo-Riemannian metrics are always of neutral signature.
Definition 9.2.
For an affine manifold $(M,\nabla)$, a 1-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian immersion of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an immersed Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the pseudo-Riemannian metric induced by $\nabla$.
The following theorem shows that space-like immersed Lagrangians are precisely those Hessian manifolds of Koszul type.
@@
Let $(M,\nabla)$ be an affine manifold, and let $\eta$ be a 1-form on $M$. Then, the following statements are equivalent:
@@
Recall that a smooth cross-section of a fiber bundle is always an immersion.
Consider canonical coordinates $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ on $T^{*}M$ such that $x^{1},\dots,x^{n}$ are affine coordinates of $M$. Now, it suffices to notice that, by the very definition, we have
$\nabla\eta=\frac{\partial\eta_{i}}{\partial x^{j}}dx^{i}\otimes dx^{j}=\eta^{*}g$
where $g$ is the pseudo-Riemannian metric induced by $\nabla$. ∎
For completeness, we recall another necessary and sufficient condition for a compact affine manifold to admit a Hessian metric of Koszul type.
@@
Combine Corollary6.26 with Theorem9.3&9.6. ∎
10 Riemannian Geometric Aspect

### v2

@@ 2.3.3 The obstruction line bundle
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type, and $(\mathscr{L},D)$ its obstruction line bundle. Then there exists a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}\rightarrow M$ such that the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is a $1$-form potential of $((M,\nabla),h)$. Moreover, if it is assumed further that there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$, then $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle.
@@
3 Interaction with Classical Geometric Structures
3.1 Digression on Complex Geometry
This digression demonstrates the utility of our framework beyond affine geometry. By lifting the Hessian structure to the tangent bundle, we can leverage powerful theorems from complex analysis.
We prove that every positive flat line bundles over an affine manifold $(M,\nabla)$ can be lifted to a positive holomorphic line bundles over the Sasaki lifting of $(M,\nabla)$, and that the Sasaki lifting of a complete Hessian manifold is Stein, which finally results in an analytic embedding theorem for complete Hessian manifolds.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon TM\rightarrow M$ be the tangent bundle of $M$. Denote by $\mathcal{A}$ the sheaf of affine functions on $(M,\nabla)$.
Consider any coordinate map
@@ 3.1 Digression on Complex Geometry
of $M$ such that $x^{1},\dots,x^{n}\in\mathcal{A}(U)$ are affine functions, and that $TM\rightarrow M$ is trivialized over $U$.
Then, it is readily seen that
@@
is a coordinate map of $TM$.
Now, the canonical almost complex structure $J_{\nabla}$ on the tangent bundle $T(TM)\rightarrow TM$ is defined by
@@
for all $i=1,\dots,n$. By the classical construction in [23], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$.
@@
Although a tangent bundle of positive dimension is never compact, we may hope that the Sasaki liftings are Stein. Fortunately, for the the Sasaki liftings of complete Hessian manifolds, this is indeed true.
Proposition 3.3.
Let $((M,\nabla),h)$ be a complete Hessian manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Then $(TM,J_{\nabla})$ is Stein.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. By considering the Sasaki lifting of $(\tilde{M},\pi^{*}\nabla)$, it is readily seen that there exists a holomorphic covering map $T\tilde{M}\rightarrow TM$. Therefore it suffices to show that $T\tilde{M}$ is Stein.
But, by Theorem2.13, $T\tilde{M}$ is a convex tube domain, and is therefore pesudoconvex by complex analysis of several variables. ∎
Now, we are ready to state and prove the main theorem of this section.
Theorem 3.4.
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper real analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
Proof.
By Proposition3.3, the Sasaki lifting $(TM,J_{\nabla})$ of $(M,\nabla)$ is Stein. By Stein’s embedding theorem [24], there exists a proper holomorphic embedding $\Phi$ of $(TM,J_{\nabla})$ into $\mathbb{C}^{d}$ for any integer $d>(3n+1)/2$. Now, it suffices to restrict the real part $\operatorname{Re}(\Phi)$ of $\Phi$ to the totally real submanifold $M$ of $(TM,J_{\nabla})$. ∎
Remark 3.5.
Recall that Whitney’s embedding theorem states that all real analytic $n$-manifolds can be analytically embedded into $\mathbb{R}^{2n}$. Theorem3.4 suggests a more efficient bound on analytical embedding dimension for complete Hessian manifolds.
3.2 Legendre Transformation
It turns out that there exists a natural duality between radiant Hessian manifolds and Hessian manifolds of Koszul type. Also, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, we prove that its dual $((M,\nabla^{*}),h)$ is also a radiant Hessian manifold of Koszul type, and provide an explicit expression of the $1$-form potential of $((M,\nabla^{*}),h)$ in terms of that of $((M,\nabla),h)$, resembling the classical Legendre transformation.
In this section, we introduce radiant manifolds and then study Hessian manifold in terms of the generalized Legendre duality mentioned above.
@@ 3.2 Legendre Transformation
It is readily seen that, if $X$ and $Y$ are vector fields on an affine manifold such that $X$ is Eulerian, then $Y$ is also Eulerian if and only if XYX-Y is parallel.
@@
By Theorem2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$.
Since, up to a translation, $\Omega$ is a cone with apex at the origin, all affine automorphisms of $(\Omega,\pi^{*}\nabla)$ are in fact linear. Now, it suffices to descend the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
defined on $(\Omega,\pi^{*}\nabla)$ to an Euler vector field of $(M,\nabla)$. ∎
The converse statement also holds if we assume compactness.
Proposition 3.11.
[…]
Proof.
By Theorem2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
@@
on $(\Omega,\pi^{*}\nabla)$, where $a^{1},\dots,a^{n}\in\mathbb{R}$. Moreover, by Poincaré-Hopf index theorem, we have $(a^{1},\dots,a^{n})\notin\Omega$. Up to an affine change of coordinates, we may assume $a^{1},\dots,a^{n}=0$. Since $M$ is compact, it follows that $H$ is a complete vector field. Since $\pi\colon\Omega\rightarrow M$ is a covering space, we obtain that $\pi^{*}H$ is also a complete vector field as $H$ is. Now it remains to recall that the integral curves of the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
@@
Corollary 3.17.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type. Then the Hessian manifold $((M,\nabla^{*}),h)$ is also radiant of Koszul type.
Proof.
This follows immediately from Proposition3.15 and Proposition3.16. ∎
Now, in view of Corollary3.17, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$. We will now give an satisfactory answer to this question.
For a vector field $X$, as per usual, we denote by $\mathcal{L}_{X}$ the Lie derivative along $X$.
Definition 3.18.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form $\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$ is said to be the Legendre transform of $\eta$ along direction $H$.
Theorem 3.19.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. Then the Legendre transform $\eta^{*}$ of $\eta$ along direction $H$ satisfies $h=\nabla^{*}\eta^{*}$.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$ so that
$H=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is Eulerian. For each $i\in\{1,\dots,n\}$ define $p_{i}:=\eta(\partial/\partial x^{i})$. Then
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}$
and $\nabla^{*}(dp_{1})=\cdots=\nabla^{*}(dp_{n})=0$. By Cartan’s magic formula, we obtain
$\eta^{*}=-\eta+\mathcal{L}_{H}\eta=d(\eta(H))-\eta=d(x^{1}p_{1}+\cdots+x^{n}p_{n})-\sum_{i=1}^{n}p_{i}dx^{i}=\sum_{i=1}^{n}x^{i}dp_{i}$
and hence
$\nabla^{*}\eta^{*}=\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=h$
as required. ∎
@@
then the construction in Definition3.18 recovers the Legendre transformation in classical mechanics.
3.3 Space-like Lagrangian Immersions
@@
Definition 3.21 (Patterson-Walker,1952[16]).
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$ with Christoffel symbol $\Gamma^{k}_{ij}$, the pseudo-Riemannian metric
$g:=\frac{1}{2}(dx^{i}\otimes dp_{i}+dp_{i}\otimes dx^{i})-\Gamma^{k}_{ij}p_{k}dx^{i}dx^{j}$
on $T^{*}M$ is called the Riemann extension of $\nabla$, where $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ here are canonical coordinates on $T^{*}M$.
Note that, the Riemann extension of a flat torsion-free connection is of neutral signature.
@@
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$, a $1$-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian immersion of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an immersed Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the Riemann extension of $\nabla$.
The following theorem shows that, in the category of affine manifolds, space-like immersed Lagrangians are precisely those Hessian manifolds of Koszul type.
@@
Let $M$ be a differentiable manifold, and let $\nabla$ be a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$. Then, for a $1$-form $\eta$ on $M$, the following statements are equivalent:
  1. 1.
    $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$;
  2. 2.
    $\eta$ is closed, and $\nabla\eta>0$.
Moreover, for a closed $1$-form $\eta$, the $(0,3)$-tensor $\nabla(\nabla\eta)$ is totally symmetric if and only if $R^{m}_{ijk}\eta_{m}=0$, where $R$ here is the curvature tensor of $\nabla$.
Proof.
Recall from differential topology that a smooth cross-section of a fiber bundle is always an immersion.
Consider canonical coordinates $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ on $T^{*}M$. Denote by $g$ the Riemann extension of $\nabla$. Now, it suffices to notice that, by the very definition, we have
$\nabla\eta=(\partial_{i}\eta_{j}-\Gamma^{k}_{ij}\eta_{k})dx^{i}\otimes dx^{j}=d\eta_{i}\otimes dx^{i}-\Gamma^{k}_{ij}\eta_{k}dx^{i}dx^{j}=\eta^{*}g$
where $\Gamma^{k}_{ij}$ is the Christoffel symbol of $\nabla$. Denote $h:=\nabla\eta$ and denote $e_{i}:=\partial/\partial x^{i}$. Then
$\nabla_{e_{k}}h=(\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik})dx^{i}\otimes dx^{j}.$
Therefore, $\nabla h$ is totally symmetric if and only if $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$. Define $S^{m}_{ijk}:=\partial_{k}\Gamma^{m}_{ij}-\Gamma^{m}_{il}\Gamma^{l}_{jk}-\Gamma^{m}_{jl}\Gamma^{l}_{ik}$. Then a straightforward computation yields that
$\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}=\partial_{i}\partial_{j}\eta_{k}-S^{m}_{ijk}\eta_{m}-(\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}).$
Notice that $\partial_{i}\partial_{j}\eta_{k}$ and $\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}$ are already invariant under permutations of indices $i,j,k$. Therefore, the expression $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$ if and only if $S^{m}_{ijk}\eta_{m}$ is. But, by construction, we have $S^{m}_{ijk}-S^{m}_{ikj}=-R^{m}_{ijk}$. This concludes the proof. ∎
Corollary 3.24.
Let $(M,\nabla)$ be an affine manifold, and $\eta$ a $1$-form on $M$. Then, the following statements are equivalent:
@@
Recall Theorem3.13 and then apply Theorem3.23. ∎
We remark that, in [26], J.Zhang and G.Khan derived a similar condition for totally symmetry of the $(0,3)$-tensor described in Theorem3.23.
3.4 Hyperbolic Affine Manifolds
For completeness, we recall yet another necessary and sufficient condition for a compact affine manifold to admit a Hessian metric of Koszul type.
@@ 3.4 Hyperbolic Affine Manifolds
Combine Corollary2.48&3.24 with Theorem3.27. ∎
We shall now introduce an important property of compact hyperbolic affine manifolds.
[…]
Definition 3.29.
Let $(M,\nabla)$ be an oriented affine manifold, and let
[…]
be the holonomy of $\nabla$. The homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}(\operatorname{det}(\rho(g)))$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$, is termed the canonical class of $(M,\nabla)$.
The canonical class of an oriented affine manifold is by no means an unfamiliar object: We now show that, when the oriented affine manifold is equipped with a complete Hessian metric, its first Koszul form represents precisely the canonical class.
Lemma 3.30.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Then, under the identification $\operatorname{H}^{1}(M;\mathbb{R})=\operatorname{H}^{1}(\pi_{1}(M);\mathbb{R})$, the cohomology class represented by the first Koszul form of $((M,\nabla),h)$ is the canonical class of $(M,\nabla)$.
Proof.
Denote by $\varkappa\in\Omega^{1}(M)$ the first Koszul form of $((M,\nabla),h)$. Since
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, it follows that $\varkappa$ is the connection 1-form of the induced connection $\nabla$ on $\bigwedge^{n}T^{*}M$. Therefore
$\operatorname{log}(\operatorname{det}(\rho(g)))=\int_{\gamma}\varkappa$
[…]
The following corollary is natural.
Corollary 3.31.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Suppose that the canonical class of $(M,\nabla)$ is a rational cohomology class. Then the canonical class of $(M,\nabla^{*})$ is also rational.
Proof.
The desired result follows immediately from Lemma3.30 and the fact that the determinant of $h$ in affine coordinates of $(M,\nabla^{*})$ is the reciprocal of the determinant of $h$ in affine coordinates of $(M,\nabla)$. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be deformed into another hyperbolic affine structure of which canonical class is rational.
To proceed, we adapt a theorem of J.-L.Koszul to our setting.
Theorem 3.32 (Koszul, 1968[17]).
[…]
Applying Theorem3.32 together with Lie algebra-valued group cohomology, we can prove the following density theorem:
Theorem 3.33.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a flat and torsion-free connection $\nabla^{\prime}$ on the tangent bundle $TM\rightarrow M$ such that $(M,\nabla^{\prime})$ is a hyperbolic affine manifold of which canonical class is rational.
Proof.
[…]
$\operatorname{Hom}(G,\operatorname{GL}^{+}_{n}(\mathbb{R}))\rightarrow\operatorname{Hom}(G,\mathbb{R}^{+}),\quad\rho\mapsto(g\mapsto\operatorname{det}(\rho(g)))$
is a submersion, or equivalently, the induced map
$\operatorname{tr}_{*}\colon\mathrm{H}^{1}(G;\mathfrak{gl}_{n}(\mathbb{R}))\rightarrow\mathrm{H}^{1}(G;\mathbb{R})$
of the trace homomorphism
$\operatorname{tr}\colon\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}$
is surjective, which is a direct consequence of the fact that the short exact sequence
$0\rightarrow\mathfrak{sl}_{n}(\mathbb{R})\rightarrow\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}\rightarrow 0$
splits. This concludes the proof. ∎
3.5 Riemannian Geometric Aspect